OpenStudy (anonymous):

What is the slope of the equation y+6=1/3(x-4) A.-1/3 B.1/3 C.4 D.6

OpenStudy (maddy1251):

What I would do is try to get the equation into the worldwide known form of y=mx+b ; where m is our slope. Do you know how to do this for this question?

OpenStudy (maddy1251):

y+6=1/3(x-4) First step; Distribute the "1/3" to the (x-4) y+6=1/3x-4/3

OpenStudy (maddy1251):

Then, subtract 6 from both sides. y+6-6 = 1/3x-4/3-6 y=1/3x-11/15

OpenStudy (maddy1251):

Now the equation is in proper y=mx+b form. \[y=\frac{ 1 }{ 3 }x - \frac{ 11 }{ 15 }\]